Hello, I would like some help with this question:
The lenght of a rectangular garden is 5 feet longer than twice its width.If the perimeter of th garden is 130 feet, ind its dimensions.

Thanks

Answer by rfer(16322)

2x+2(2x+5)=130
2x+4x+10=130
6x=120
x=120/6
x=20
2x+5=45